# Semantic Agreement Enables Efficient Open-Ended LLM Cascades

### v1

@@ H.3 Latencies at 98% Target Model Performance
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 1594 1541 1509 0.1
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 1608 1545 1533 0.1
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 1745 1659 1606 0.1
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 1662 1622 1538 0.1
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 1776 1682 1629 0.1
Llama3.1-8B, Qwen2.5-7B 1661 1477 1442 0.1
Llama3.1-8B, Qwen2.5-3B 1901 1681 1677 0.2
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 1682 1463 1412 0.1
Llama3.1-8B, Qwen2.5-7B, mT0-Large 1735 1477 1454 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
1564 1594 1571 1568 1309 1538
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
1719 1696 1706 1705 1331 1666
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
1660 1645 1661 1670 1359 1617
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
1725 1770 1749 1774 1503 1713
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
1620 1661 1625 1642 1330 1639
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
1715 1727 1735 1744 1339 1691
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
1447 1575 1482 1586 1025 1460
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
1475 1695 1536 1666 861.4 1597
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
1580 1550 1592 1572 1036 1530
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
1507 1456 1434 1465 962.2 1375
@@ H.3.4 CNN/DailyMail
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 7824 5787 5123 4705 0.2
Qwen2.5-3B 7217 5982 5803 5318 0.5
Qwen2.5-1.5B 7095 6790 5196 4639 1.0
Qwen2.5-0.5B 6682 6609 6098 5752 0.9
Gemma3-1B 6954 6264 $59$ 5779 0.3
Llama3.2-1B 6890 5808 5489 5356 0.2
Llama3.2-3B 7585 6251 6284 5574 0.1
Llama3.1-8B 8936 7947 7701 6818 0.0
mT0-Large 6964 6062 5677 5444 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Llama3.2-1B, mT0-Large 5511 5650 5172 0.1
Llama3.2-1B, Qwen2.5-1.5B 5714 6318 5462 0.2
Qwen2.5-1.5B, Gemma3-1B 6670 5827 4626 1.0
Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 5103 6564 5402 0.2
Qwen2.5-7B, Llama3.1-8B 7011 6692 5763 0.2
Qwen2.5-7B, mT0-Large 5733 6291 5806 0.4
Qwen2.5-7B, Qwen2.5-3B 5541 5621 4751 0.4
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Llama3.2-1B, mT0-Large,
Qwen2.5-1.5B
3816 3245 4174 4055 5070 4187
Llama3.2-1B, mT0-Large,
Qwen2.5-0.5B
5039 4873 4986 4999 6254 5570
Qwen2.5-1.5B, Qwen2.5-0.5B,
Gemma3-1B
5422 5834 6551 6444 6199 6571
Llama3.2-1B, Gemma3-1B,
Qwen2.5-0.5B
5262 5149 5235 4790 5799 5096
Llama3.2-3B, Qwen2.5-3B,
Qwen2.5-0.5B
5733 5242 5408 5448 6444 5322
Qwen2.5-7B, mT0-Large,
Llama3.2-3B
4041 4744 4691 4997 6125 4525
Qwen2.5-7B, Llama3.1-8B,
mT0-Large
5059 6022 5557 5431 7310 5803
Llama3.1-8B, mT0-Large,
Qwen2.5-3B
6633 7257 6945 6925 $75$ 6745
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 557.1 215 221.3 212.3 0.1
Qwen2.5-3B 464.1 302.8 298.4 297 0.1
Qwen2.5-1.5B 442.5 281.2 300.6 296.2 0.1
Qwen2.5-0.5B 419.5 365.1 364.3 368.9 0.1
Gemma3-1B 435.2 411.4 409.2 405.1 0.4
Llama3.2-1B 419.2 385.2 375.4 372.7 0.3
Llama3.2-3B 455.3 393.8 373 310.6 0.5
Llama3.1-8B 536.9 247.1 273.9 247.1 0.0
FLAN-T5-Large 453.3 314.7 432.5 348.8 0.0
mT0-Large 442.1 130.7 118.7 109.1 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
mT0-Large, Qwen2.5-1.5B 135.9 146 184.6 0.0
mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large 339.3 227.2 231.8 0.0
Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B 398.1 392 389.6 0.0
Qwen2.5-7B, Llama3.1-8B 212.3 224.3 203.6 0.0
Qwen2.5-7B, Llama3.1-8B, mT0-Large 163.1 163.1 163.1 0.0
Qwen2.5-7B, Llama3.2-3B 467.4 455.9 415.2 0.5
Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 513 540.7 450.7 0.5
Qwen2.5-7B, mT0-Large, FLAN-T5-Large 441.7 171.3 230.9 0.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
mT0-Large, Qwen2.5-1.5B,
FLAN-T5-Large
122.2 98.95 217.9 98.95 149.8 86.65
mT0-Large, Qwen2.5-1.5B,
FLAN-T5-Large, Llama3.2-1B
152 122.2 214.6 117.5 179.1 159.9
Qwen2.5-1.5B, FLAN-T5-Large,
Gemma3-1B
291.4 285.4 329.2 285.1 291.4 316.6
Qwen2.5-1.5B, FLAN-T5-Large,
Gemma3-1B, Llama3.2-1B
327.2 314.1 314.4 314.4 289.2 335.4
Qwen2.5-1.5B, FLAN-T5-Large,
Llama3.2-1B
318.5 302.1 313.6 301.5 281.9 293.3
Llama3.2-3B, Qwen2.5-3B,
FLAN-T5-Large
216.4 232 250.8 236.1 176.2 239.1
Qwen2.5-7B, mT0-Large,
FLAN-T5-Large
163.1 163.1 213.9 163.1 163.1 163.1
Qwen2.5-7B, mT0-Large,
Llama3.2-3B
176.5 163.1 231.7 163.1 163.1 163.1
Qwen2.5-7B, Llama3.1-8B,
mT0-Large
163.1 163.1 200.8 163.1 163.1 163.1
Llama3.1-8B, mT0-Large,
Qwen2.5-1.5B
173.9 168.1 277.2 169.8 168.4 156.1
@@ H.3.6 TriviaQA
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 448.7 332.8 333.8 334.2 0.1
Qwen2.5-3B 402.7 335.9 333.4 335.5 0.1
Qwen2.5-1.5B 392.6 351.4 351.1 351.7 0.0
Qwen2.5-0.5B 386.2 377.4 378.8 378.2 0.0
Gemma3-1B 385.1 373 374.9 373.3 0.0
Llama3.2-1B 379.1 353.9 353.8 353.9 0.0
Llama3.2-3B 405.8 302.5 303.2 301.6 0.1
Llama3.1-8B 490.2 311.2 328 336.5 0.1
FLAN-T5-Large 388.8 372.3 379 371.4 0.0
mT0-Large 400.8 396.4 396.6 396.8 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-1.5B, Llama3.2-1B 356.8 358.4 357.7 0.0
Llama3.2-1B, Gemma3-1B 361.2 361.9 360.5 0.0
Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 358.8 359.3 356.8 0.1
Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 366.5 367.3 368.5 0.1
Llama3.1-8B, Qwen2.5-7B 359.3 390.7 354.9 0.1
Llama3.1-8B, Qwen2.5-7B, mT0-Large 490.3 492.4 491.9 0.0
Qwen2.5-7B, Llama3.2-3B 332.1 331.7 331.3 0.0
Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 404.3 429.9 404 0.5
Qwen2.5-7B, Gemma3-1B 433.3 434.5 433.7 0.0
Llama3.2-3B, Qwen2.5-3B 321.7 331.5 325.6 0.1
Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 400.6 415.3 394.9 0.5
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Qwen2.5-1.5B, Llama3.2-1B,
Gemma3-1B
360.6 354.7 378.3 355.4 357.2 361.5
Qwen2.5-1.5B, Llama3.2-1B,
FLAN-T5-Large
367.9 364.3 381.6 364.7 364.5 370.9
Llama3.2-1B, Gemma3-1B,
FLAN-T5-Large
368.7 364.4 380.7 364.3 364.4 369
Qwen2.5-1.5B, Llama3.2-1B,
Gemma3-1B, FLAN-T5-Large
364.2 358.5 379.7 361.2 358.6 365.5
Llama3.1-8B, Qwen2.5-7B,
Llama3.2-3B
329.5 314.7 397.9 315.2 313.6 338.7
Qwen2.5-7B, Llama3.2-3B,
Gemma3-1B
354.7 357.9 391.9 350.4 356 370
Qwen2.5-7B, Llama3.2-3B,
Qwen2.5-3B
328.4 320.9 390.2 320.8 323.5 338.8
Llama3.1-8B, Llama3.2-3B,
Qwen2.5-3B
346.1 334.3 403.5 334.2 337.1 362.5
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
364.1 335 412.4 331.5 375.4 414.7
@@ H.4 Performances at 40% Target Model FLOPs
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.6906 0.6968 0.6948 0.2
Qwen2.5-3B 0.6667 0.6735 0.6743 0.3
Qwen2.5-1.5B 0.6506 0.6594 0.6583 0.3
Qwen2.5-0.5B 0.5483 0.5588 0.5585 0.5
Gemma3-1B 0.6441 0.6543 0.6537 0.3
Llama3.2-1B 0.5954 0.603 0.6036 0.2
Llama3.2-3B 0.6707 0.6773 0.6776 0.3
Llama3.1-8B 0.6985 0.7048 0.7043 0.2
FLAN-T5-Large 0.5584 0.5638 0.5643 0.3
mT0-Large 0.5799 0.588 0.5867 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 0.6458 0.6486 0.6527 0.1
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6413 0.6446 0.6469 0.1
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.641 0.6458 0.6503 0.2
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.6424 0.6448 0.65 0.1
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6456 0.6436 0.6457 0.3
Llama3.1-8B, Qwen2.5-7B 0.6924 0.6956 0.6958 0.0
Llama3.1-8B, Qwen2.5-3B 0.6871 0.6889 0.6876 0.2
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6903 0.6929 0.6943 0.0
Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6913 0.6942 0.6927 0.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
0.6468 0.6487 0.6478 0.6507 0.6712 0.6545
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
0.6473 0.6521 0.6512 0.6535 0.6801 0.6576
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
0.6467 0.6491 0.6496 0.6503 0.6796 0.6595
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
0.6296 0.6339 0.6298 0.6331 0.6649 0.6413
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
0.6436 0.6483 0.6475 0.6498 0.6786 0.6545
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
0.6442 0.6472 0.6482 0.6501 0.682 0.6571
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
0.6934 0.6947 0.6942 0.6947 0.7081 0.6965
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
0.6944 0.6956 0.6958 0.695 0.7112 0.6963
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
0.6932 0.6946 0.6955 0.6944 0.7079 0.6957
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
0.6676 0.6719 0.6703 0.6729 0.6937 0.6779
@@
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.7457 0.745 0.7454 0.0
Qwen2.5-3B 0.737 0.7381 0.7388 0.1
Qwen2.5-1.5B 0.7367 0.7374 0.738 0.1
Qwen2.5-0.5B 0.7127 0.7188 0.7181 0.2
Gemma3-1B 0.7268 0.7291 0.7301 0.8
Llama3.2-1B 0.7078 0.7133 0.7134 0.1
Llama3.2-3B 0.7381 0.7393 0.7391 0.7
Llama3.1-8B 0.743 0.7441 0.7445 0.1
FLAN-T5-Large 0.7169 0.7245 0.7244 0.1
mT0-Large 0.7023 0.7099 0.7091 0.4
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 0.7334 0.7333 0.737 0.5
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.7263 0.7252 0.7259 0.1
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.725 0.726 0.7276 0.2
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.7317 0.7319 0.7314 0.3
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.7225 0.7197 0.7222 0.0
Llama3.1-8B, Qwen2.5-7B 0.7446 0.7449 0.7449 0.2
Llama3.1-8B, Qwen2.5-3B 0.7414 0.7411 0.7433 0.7
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.741 0.7371 0.7391 0.1
Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.7343 0.731 0.7416 1.0

### v3

@@ H.3 Latencies at 98% Target Model Performance
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 1594 1541 1509 0.1 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 1608 1545 1533 0.1 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 1745 1659 1606 0.1 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 1662 1622 1538 0.1 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 1776 1682 1629 0.1 Llama3.1-8B, Qwen2.5-7B 1661 1477 1442 0.1 Llama3.1-8B, Qwen2.5-3B 1901 1681 1677 0.2 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 1682 1463 1412 0.1 Llama3.1-8B, Qwen2.5-7B, mT0-Large 1735 1477 1454 0.1
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 1564 1594 1571 1568 1309 1538 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 1719 1696 1706 1705 1331 1666 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 1660 1645 1661 1670 1359 1617 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 1725 1770 1749 1774 1503 1713 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 1620 1661 1625 1642 1330 1639 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 1715 1727 1735 1744 1339 1691 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 1447 1575 1482 1586 1025 1460 Llama3.1-8B, Qwen2.5-7B, mT0-Large 1475 1695 1536 1666 861.4 1597 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 1580 1550 1592 1572 1036 1530 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 1507 1456 1434 1465 962.2 1375
@@ H.3.4 CNN/DailyMail
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 7824 5787 5123 4705 0.2 Qwen2.5-3B 7217 5982 5803 5318 0.5 Qwen2.5-1.5B 7095 6790 5196 4639 1.0 Qwen2.5-0.5B 6682 6609 6098 5752 0.9 Gemma3-1B 6954 6264 5959 5779 0.3 Llama3.2-1B 6890 5808 5489 5356 0.2 Llama3.2-3B 7585 6251 6284 5574 0.1 Llama3.1-8B 8936 7947 7701 6818 0.0 mT0-Large 6964 6062 5677 5444 0.2
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Llama3.2-1B, mT0-Large 5511 5650 5172 0.1 Llama3.2-1B, Qwen2.5-1.5B 5714 6318 5462 0.2 Qwen2.5-1.5B, Gemma3-1B 6670 5827 4626 1.0 Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 5103 6564 5402 0.2 Qwen2.5-7B, Llama3.1-8B 7011 6692 5763 0.2 Qwen2.5-7B, mT0-Large 5733 6291 5806 0.4 Qwen2.5-7B, Qwen2.5-3B 5541 5621 4751 0.4
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Llama3.2-1B, mT0-Large, Qwen2.5-1.5B 3816 3245 4174 4055 5070 4187 Llama3.2-1B, mT0-Large, Qwen2.5-0.5B 5039 4873 4986 4999 6254 5570 Qwen2.5-1.5B, Qwen2.5-0.5B, Gemma3-1B 5422 5834 6551 6444 6199 6571 Llama3.2-1B, Gemma3-1B, Qwen2.5-0.5B 5262 5149 5235 4790 5799 5096 Llama3.2-3B, Qwen2.5-3B, Qwen2.5-0.5B 5733 5242 5408 5448 6444 5322 Qwen2.5-7B, mT0-Large, Llama3.2-3B 4041 4744 4691 4997 6125 4525 Qwen2.5-7B, Llama3.1-8B, mT0-Large 5059 6022 5557 5431 7310 5803 Llama3.1-8B, mT0-Large, Qwen2.5-3B 6633 7257 6945 6925 7575 6745
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 557.1 215 221.3 212.3 0.1 Qwen2.5-3B 464.1 302.8 298.4 297 0.1 Qwen2.5-1.5B 442.5 281.2 300.6 296.2 0.1 Qwen2.5-0.5B 419.5 365.1 364.3 368.9 0.1 Gemma3-1B 435.2 411.4 409.2 405.1 0.4 Llama3.2-1B 419.2 385.2 375.4 372.7 0.3 Llama3.2-3B 455.3 393.8 373 310.6 0.5 Llama3.1-8B 536.9 247.1 273.9 247.1 0.0 FLAN-T5-Large 453.3 314.7 432.5 348.8 0.0 mT0-Large 442.1 130.7 118.7 109.1 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile mT0-Large, Qwen2.5-1.5B 135.9 146 184.6 0.0 mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large 339.3 227.2 231.8 0.0 Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B 398.1 392 389.6 0.0 Qwen2.5-7B, Llama3.1-8B 212.3 224.3 203.6 0.0 Qwen2.5-7B, Llama3.1-8B, mT0-Large 163.1 163.1 163.1 0.0 Qwen2.5-7B, Llama3.2-3B 467.4 455.9 415.2 0.5 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 513 540.7 450.7 0.5 Qwen2.5-7B, mT0-Large, FLAN-T5-Large 441.7 171.3 230.9 0.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large 122.2 98.95 217.9 98.95 149.8 86.65 mT0-Large, Qwen2.5-1.5B, FLAN-T5-Large, Llama3.2-1B 152 122.2 214.6 117.5 179.1 159.9 Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B 291.4 285.4 329.2 285.1 291.4 316.6 Qwen2.5-1.5B, FLAN-T5-Large, Gemma3-1B, Llama3.2-1B 327.2 314.1 314.4 314.4 289.2 335.4 Qwen2.5-1.5B, FLAN-T5-Large, Llama3.2-1B 318.5 302.1 313.6 301.5 281.9 293.3 Llama3.2-3B, Qwen2.5-3B, FLAN-T5-Large 216.4 232 250.8 236.1 176.2 239.1 Qwen2.5-7B, mT0-Large, FLAN-T5-Large 163.1 163.1 213.9 163.1 163.1 163.1 Qwen2.5-7B, mT0-Large, Llama3.2-3B 176.5 163.1 231.7 163.1 163.1 163.1 Qwen2.5-7B, Llama3.1-8B, mT0-Large 163.1 163.1 200.8 163.1 163.1 163.1 Llama3.1-8B, mT0-Large, Qwen2.5-1.5B 173.9 168.1 277.2 169.8 168.4 156.1
@@ H.3.6 TriviaQA
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 448.7 332.8 333.8 334.2 0.1 Qwen2.5-3B 402.7 335.9 333.4 335.5 0.1 Qwen2.5-1.5B 392.6 351.4 351.1 351.7 0.0 Qwen2.5-0.5B 386.2 377.4 378.8 378.2 0.0 Gemma3-1B 385.1 373 374.9 373.3 0.0 Llama3.2-1B 379.1 353.9 353.8 353.9 0.0 Llama3.2-3B 405.8 302.5 303.2 301.6 0.1 Llama3.1-8B 490.2 311.2 328 336.5 0.1 FLAN-T5-Large 388.8 372.3 379 371.4 0.0 mT0-Large 400.8 396.4 396.6 396.8 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-1.5B, Llama3.2-1B 356.8 358.4 357.7 0.0 Llama3.2-1B, Gemma3-1B 361.2 361.9 360.5 0.0 Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 358.8 359.3 356.8 0.1 Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 366.5 367.3 368.5 0.1 Llama3.1-8B, Qwen2.5-7B 359.3 390.7 354.9 0.1 Llama3.1-8B, Qwen2.5-7B, mT0-Large 490.3 492.4 491.9 0.0 Qwen2.5-7B, Llama3.2-3B 332.1 331.7 331.3 0.0 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 404.3 429.9 404 0.5 Qwen2.5-7B, Gemma3-1B 433.3 434.5 433.7 0.0 Llama3.2-3B, Qwen2.5-3B 321.7 331.5 325.6 0.1 Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 400.6 415.3 394.9 0.5
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B 360.6 354.7 378.3 355.4 357.2 361.5 Qwen2.5-1.5B, Llama3.2-1B, FLAN-T5-Large 367.9 364.3 381.6 364.7 364.5 370.9 Llama3.2-1B, Gemma3-1B, FLAN-T5-Large 368.7 364.4 380.7 364.3 364.4 369 Qwen2.5-1.5B, Llama3.2-1B, Gemma3-1B, FLAN-T5-Large 364.2 358.5 379.7 361.2 358.6 365.5 Llama3.1-8B, Qwen2.5-7B, Llama3.2-3B 329.5 314.7 397.9 315.2 313.6 338.7 Qwen2.5-7B, Llama3.2-3B, Gemma3-1B 354.7 357.9 391.9 350.4 356 370 Qwen2.5-7B, Llama3.2-3B, Qwen2.5-3B 328.4 320.9 390.2 320.8 323.5 338.8 Llama3.1-8B, Llama3.2-3B, Qwen2.5-3B 346.1 334.3 403.5 334.2 337.1 362.5 Llama3.1-8B, Qwen2.5-7B, mT0-Large 364.1 335 412.4 331.5 375.4 414.7
@@ H.4 Performances at 40% Target Model FLOPs
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.6906 0.6968 0.6948 0.2 Qwen2.5-3B 0.6667 0.6735 0.6743 0.3 Qwen2.5-1.5B 0.6506 0.6594 0.6583 0.3 Qwen2.5-0.5B 0.5483 0.5588 0.5585 0.5 Gemma3-1B 0.6441 0.6543 0.6537 0.3 Llama3.2-1B 0.5954 0.603 0.6036 0.2 Llama3.2-3B 0.6707 0.6773 0.6776 0.3 Llama3.1-8B 0.6985 0.7048 0.7043 0.2 FLAN-T5-Large 0.5584 0.5638 0.5643 0.3 mT0-Large 0.5799 0.588 0.5867 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 0.6458 0.6486 0.6527 0.1 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6413 0.6446 0.6469 0.1 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.641 0.6458 0.6503 0.2 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.6424 0.6448 0.65 0.1 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6456 0.6436 0.6457 0.3 Llama3.1-8B, Qwen2.5-7B 0.6924 0.6956 0.6958 0.0 Llama3.1-8B, Qwen2.5-3B 0.6871 0.6889 0.6876 0.2 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6903 0.6929 0.6943 0.0 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6913 0.6942 0.6927 0.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6468 0.6487 0.6478 0.6507 0.6712 0.6545 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6473 0.6521 0.6512 0.6535 0.6801 0.6576 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 0.6467 0.6491 0.6496 0.6503 0.6796 0.6595 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6296 0.6339 0.6298 0.6331 0.6649 0.6413 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6436 0.6483 0.6475 0.6498 0.6786 0.6545 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 0.6442 0.6472 0.6482 0.6501 0.682 0.6571 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6934 0.6947 0.6942 0.6947 0.7081 0.6965 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6944 0.6956 0.6958 0.695 0.7112 0.6963 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 0.6932 0.6946 0.6955 0.6944 0.7079 0.6957 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 0.6676 0.6719 0.6703 0.6729 0.6937 0.6779
@@
Model Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.7457 0.745 0.7454 0.0 Qwen2.5-3B 0.737 0.7381 0.7388 0.1 Qwen2.5-1.5B 0.7367 0.7374 0.738 0.1 Qwen2.5-0.5B 0.7127 0.7188 0.7181 0.2 Gemma3-1B 0.7268 0.7291 0.7301 0.8 Llama3.2-1B 0.7078 0.7133 0.7134 0.1 Llama3.2-3B 0.7381 0.7393 0.7391 0.7 Llama3.1-8B 0.743 0.7441 0.7445 0.1 FLAN-T5-Large 0.7169 0.7245 0.7244 0.1 mT0-Large 0.7023 0.7099 0.7091 0.4
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 0.7334 0.7333 0.737 0.5 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.7263 0.7252 0.7259 0.1 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.725 0.726 0.7276 0.2 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.7317 0.7319 0.7314 0.3 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.7225 0.7197 0.7222 0.0 Llama3.1-8B, Qwen2.5-7B 0.7446 0.7449 0.7449 0.2 Llama3.1-8B, Qwen2.5-3B 0.7414 0.7411 0.7433 0.7 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.741 0.7371 0.7391 0.1 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.7343 0.731 0.7416 1.0